# Efficient designs for multivariate crossover trials

### v1

@@ Appendix B Efficiency under the generalized markov-type structure
$\begin{split}\bm{\tilde{C}}_{d(s2)}&=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)\\ &=\bm{\tilde{C}}_{d(s2)(11)}-\bm{\tilde{C}}_{d(s2)(12)}\bm{\tilde{C}}^{-}_{d(s2)(22)}\bm{\tilde{C}}_{d(s2)(21)},\end{split}$ (S.76)
where $\bm{\tilde{C}}_{d(s2)(11)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$,
$\bm{\tilde{C}}_{d(s2)(12)}=\bm{\tilde{C}}^{{}^{\prime}}_{d(s2)(21)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, and
$\bm{\tilde{C}}_{d(s2)(22)}=\left(\bm{I}_{2}\otimes\bm{H}_{t}\bm{F}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$. Here, $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}$.
@@
Following the same approach as to obtain the information matrix for the direct effects in the presence of the period effects as disucssed in Lemma S.3, we can prove (S.76). ∎
@@
The matrix $\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}$ can be expressed as
$\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{\varOmega}_{1}&-\bm{I}_{n}\otimes\bm{\varOmega}_{2}\\ -\bm{I}_{n}\otimes\bm{\varOmega}_{2}&\bm{I}_{n}\otimes\bm{\varOmega}_{4}\end{bmatrix},$ (S.80)
where $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$, and $\bm{\varOmega}_{4}$ are as given in Lemma S.4.
@@
$\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)$ $=\begin{bmatrix}\bm{A}^{(5)}&\bm{B}^{(5)}\\ \bm{B}^{(5)}&\bm{D}^{(5)}\end{bmatrix},$ (S.83)
where $\bm{A}^{(5)}=\delta_{1}^{-1}\bm{I}_{n}$, $\bm{B}^{(5)}=-\bar{\rho}\delta_{2}^{-1}\bm{I}_{n}$, and $\bm{D}^{(5)}=\delta_{2}^{-1}\bm{I}_{n}$. Here, $\delta_{1}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}\bm{{1}}_{p}$, $\delta_{2}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\bm{{1}}_{p}$, $\bm{K}_{1}^{-1}=\bm{V}_{1}^{-1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}$ and $\bm{K}_{2}=\sigma_{12}\bm{V}_{R}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$. So,
$\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}=\\ \begin{bmatrix}\left(\bm{A}^{(5)}\right)^{-}+\left(\bm{A}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{Z}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{A}^{(5)}\right)^{-}&-\left(\bm{A}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{Z}^{(5)}\right)^{-}\\ -\left(\bm{Z}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{A}^{(5)}\right)^{-}&\left(\bm{Z}^{(5)}\right)^{-}\end{bmatrix},$ (S.84)
@@
where $\delta_{4}=\delta_{2}^{-1}-\bar{\rho}^{2}\delta_{1}\delta_{2}^{-2}$. Using the expression $\delta_{4}$, we get that $\delta_{1}+\bar{\rho}^{2}\delta_{4}^{-1}\delta_{1}^{2}\delta_{2}^{-2}=\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}$. So, from (S.84) and (S.85), we get
$\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}$ $=\delta_{4}^{-1}\begin{bmatrix}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}&\bar{\rho}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}\\ \bar{\rho}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}&\bm{I}_{n}\end{bmatrix}.$ (S.88)
@@
$\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}&-\bm{I}_{n}\otimes\bar{\rho}\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\\ -\bm{I}_{n}\otimes\bar{\rho}\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}&\bm{I}_{n}\otimes\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\end{bmatrix}.$ (S.91)
@@
$\bm{\varSigma}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\varSigma}^{-1}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{A}^{(6)}&-\bm{I}_{n}\otimes\bm{B}^{(6)}\\ -\bm{I}_{n}\otimes\bm{B}^{(6)}&\bm{I}_{n}\otimes\bm{D}^{(6)}\end{bmatrix},$ (S.94)
@@
Here $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$. Thus from (S.30), (S.31), (S.51), (S.52) and (S.94), by calculation we can prove (S.80). ∎
@@
For any design $d\in\varOmega_{t,n,p}$, $\bm{{C}}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$. If $d$ is a design uniform on periods, then $\bm{{C}}_{d(s2)}=\bm{\tilde{C}}_{d(s2)}$.
@@
Since $pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{Z}_{1}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{P}&\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)\leq pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)$, where $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}$, using the expression of the matrices $\bm{C}_{d(s2)}$ and $\bm{\tilde{C}}_{d(s2)}$ as given in Lemma S.3 and Lemma S.5, respectively, we get that $\bm{C}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$.
@@
Since $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, where $\bm{\psi}=\begin{bmatrix}\bm{{0}}^{{}^{\prime}}_{p-1\times 1}&0\\ \bm{I}_{p-1}&\bm{{0}}_{p-1\times 1}\end{bmatrix}$, using the expression of the matrices $\bm{\tilde{C}}_{d(s2)(11)}$, $\bm{\tilde{C}}_{d(s2)(12)}$, $\bm{\tilde{C}}_{d(s2)(21)}$ and $\bm{\tilde{C}}_{d(s2)(22)}$ as given in Lemma S.5, and the expression of $\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}$ as given in Lemma S.6, we get
@@
$\begin{split}\bm{\tilde{C}}_{d(s2)(11)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{\tilde{C}}_{d(s2)(12)}&=\bm{\tilde{C}}^{{}^{\prime}}_{d(s2)(21)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{\tilde{C}}_{d(s2)(22)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\end{split}$ (S.95)
where $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$, and $\bm{\varOmega}_{4}$ are the matrices as given in Lemma S.4.
Since $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, from the expression of the matrices $\bm{C}_{{d(s2)}(11)(1)}$, $\bm{C}_{{d(s2)}(12)(1)}$, $\bm{C}_{{d(s2)}(21)(1)}$ and $\bm{C}_{{d(s2)}(22)(1)}$ as given in Lemma S.3, and using the expression of $\bm{A}^{*}$ as given in Lemma S.4, we get

### v2

@@ Appendix B Efficiency under the generalized Markov-type structure
$\begin{split}\bm{\tilde{C}}_{d(s2)}&=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)\\ &=\bm{\tilde{C}}_{d(s2)(11)}-\bm{\tilde{C}}_{d(s2)(12)}\bm{\tilde{C}}^{-}_{d(s2)(22)}\bm{\tilde{C}}_{d(s2)(21)},\end{split}$ (S.40)
where $\bm{\tilde{C}}_{d(s2)(11)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$,
$\bm{\tilde{C}}_{d(s2)(12)}=\bm{\tilde{C}}^{{}^{\prime}}_{d(s2)(21)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$, and
$\bm{\tilde{C}}_{d(s2)(22)}=\left(\bm{I}_{2}\otimes\bm{H}_{t}\bm{F}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$.
@@
Following the same approach as to obtain the information matrix for the direct effects in the presence of the period effects as discussed in Lemma S.3, we can prove (S.40). ∎
@@
The matrix $\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}$ can be expressed as
$\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}&-\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\\ -\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}&\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\end{bmatrix},$ (S.41)
where $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$, and $\bm{\mathit{\Omega}}_{4}$ are as given in Lemma S.4.
@@
$\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)$ $=\begin{bmatrix}\bm{A}^{(5)}&\bm{B}^{(5)}\\ \bm{B}^{(5)}&\bm{D}^{(5)}\end{bmatrix},$ (S.42)
where $\bm{A}^{(5)}=\delta_{1}^{-1}\bm{I}_{n}$, $\bm{B}^{(5)}=-\bar{\rho}\delta_{2}^{-1}\bm{I}_{n}$, and $\bm{D}^{(5)}=\delta_{2}^{-1}\bm{I}_{n}$. Here, $\delta_{1}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}\bm{{1}}_{p}$, $\delta_{2}=1/\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\bm{{1}}_{p}$, $\bm{K}_{1}^{-1}=\bm{V}_{1}^{-1}+\bar{\rho}^{2}\bm{K}_{2}^{-1}$ and $\bm{K}_{2}=\sigma_{12}\bm{V}_{R}$, $\bar{\rho}=\rho\sqrt{\frac{\sigma_{22}}{\sigma_{11}}}$ and $\sigma_{12}=\sigma_{22}\left(1-\rho^{2}\right)$. So
$\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}=\\ \begin{bmatrix}\left(\bm{A}^{(5)}\right)^{-}+\left(\bm{A}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{Z}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{A}^{(5)}\right)^{-}&-\left(\bm{A}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{Z}^{(5)}\right)^{-}\\ -\left(\bm{Z}^{(5)}\right)^{-}\bm{B}^{(5)}\left(\bm{A}^{(5)}\right)^{-}&\left(\bm{Z}^{(5)}\right)^{-}\end{bmatrix},$ (S.43)
@@
where $\delta_{4}=\delta_{2}^{-1}-\bar{\rho}^{2}\delta_{1}\delta_{2}^{-2}$. Using the expression $\delta_{4}$, we get that $\delta_{1}+\bar{\rho}^{2}\delta_{4}^{-1}\delta_{1}^{2}\delta_{2}^{-2}=$
$\delta_{4}^{-1}\delta_{1}\delta_{2}^{-1}$. So from (S.43) and (S.44), we get
$\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}$ $=\delta_{4}^{-1}\begin{bmatrix}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}&\bar{\rho}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}\\ \bar{\rho}\delta_{1}\delta_{2}^{-1}\bm{I}_{n}&\bm{I}_{n}\end{bmatrix}.$ (S.45)
@@
$\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{1}^{-1}&-\bm{I}_{n}\otimes\bar{\rho}\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\\ -\bm{I}_{n}\otimes\bar{\rho}\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}&\bm{I}_{n}\otimes\bm{{1}}_{p}^{{}^{\prime}}\bm{K}_{2}^{-1}\end{bmatrix}.$ (S.46)
@@
$\bm{\mathit{\Sigma}}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\left[\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}\left(\bm{I}_{2}\otimes\bm{U}\right)\right]^{-}\left(\bm{I}_{2}\otimes\bm{U}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1}$ $=\begin{bmatrix}\bm{I}_{n}\otimes\bm{A}^{(6)}&-\bm{I}_{n}\otimes\bm{B}^{(6)}\\ -\bm{I}_{n}\otimes\bm{B}^{(6)}&\bm{I}_{n}\otimes\bm{D}^{(6)}\end{bmatrix},$ (S.47)
@@
Thus from (S.18), (S.19), (S.27), (S.28) and (S.47), by calculation we can prove (S.41). ∎
@@
For any design $d\in\Omega_{t,n,p}$, $\bm{{C}}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$. If $d$ is a design uniform on periods, then $\bm{{C}}_{d(s2)}=\bm{\tilde{C}}_{d(s2)}$.
@@
Since $pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{Z_{1}}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{P}&\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)$
@@
$\leq pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)$, using the expression of the matrices $\bm{C}_{d(s2)}$ and $\bm{\tilde{C}}_{d(s2)}$ as given in Lemma S.3 and Lemma S.5, respectively, we get that $\bm{C}_{d(s2)}\leq\bm{\tilde{C}}_{d(s2)}$.
@@
Since $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, where $\bm{\psi}=\begin{bmatrix}\bm{{0}}^{{}^{\prime}}_{p-1\times 1}&0\\ \bm{I}_{p-1}&\bm{{0}}_{p-1\times 1}\end{bmatrix}$, using the expression of the matrices $\bm{\tilde{C}}_{d(s2)(11)}$, $\bm{\tilde{C}}_{d(s2)(12)}$, $\bm{\tilde{C}}_{d(s2)(21)}$ and $\bm{\tilde{C}}_{d(s2)(22)}$ as given in Lemma S.5, and the expression of $\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{2}\otimes\bm{U}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}$ as given in Lemma S.6, we get
@@
$\begin{split}\bm{\tilde{C}}_{d(s2)(11)}&=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d}\end{bmatrix},\\ \bm{\tilde{C}}_{d(s2)(12)}&=\bm{\tilde{C}}^{{}^{\prime}}_{d(s2)(21)}=\begin{bmatrix}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\text{ and}\\ \bm{\tilde{C}}_{d(s2)(22)}&=\begin{bmatrix}\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&-\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\\ -\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}&\bm{H}_{t}\bm{T}_{d}^{{}^{\prime}}\left(\bm{I}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d}\bm{H}_{t}\end{bmatrix},\end{split}$ (S.48)
where $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$, and $\bm{\mathit{\Omega}}_{4}$ are the matrices as given in Lemma S.4.
Since $\bm{F}_{d}=\left(\bm{I}_{n}\otimes\bm{\psi}\right)\bm{T}_{d}$, from the expression of the matrices $\bm{C}_{{d(s2)}(11)(1)}$, $\bm{C}_{{d(s2)}(12)(1)}$,
$\bm{C}_{{d(s2)}(21)(1)}$ and $\bm{C}_{{d(s2)}(22)(1)}$ as given in Lemma S.3, and using the expression of $\bm{A}^{*}$ as given in Lemma S.4, we get